# Self-supervised pre-training with diffusion model for few-shot landmark detection in x-ray images

### v1

@@ Abstract
In the last few years, deep neural networks have been extensively applied in the medical domain for different tasks, ranging from image classification and segmentation to landmark detection. However, the application of these technologies in the medical domain is often hindered by data scarcity, both in terms of available annotations and images. This study introduces a new self-supervised pre-training protocol based on diffusion models for landmark detection in x-ray images. Our results show that the proposed self-supervised framework can provide accurate landmark detection with a minimal number of available annotated training images (up to $50$), outperforming ImageNet supervised pre-training and state-of-the-art self-supervised pre-trainings for three popular x-ray benchmark datasets. To our knowledge, this is the first exploration of diffusion models for self-supervised learning in landmark detection, which may offer a valuable pre-training approach in few-shot regimes, for mitigating data scarcity.
@@ 1 Introduction
Landmark detection is the task of identifying anatomical keypoints of interest in images [1]. Detecting key points can directly assist in angle measurements [2], skeletal measurements [3], and surgical planning [4].
@@
In recent years, deep neural networks have become extensively utilized for the landmark detection task [5, 6], and many fully supervised methods have been proposed [7, 8, 9, 10, 11, 12, 13, 14, 15]. However, the limited availability of annotated data and the reliance on medical experts for annotations present challenges in applying deep learning to this domain. Specifically, annotating landmarks is labor-intensive, costly, and demands expert precision [8, 7, 9], particularly for critical tasks such as surgical treatment planning. Consequently, real-world datasets typically contain very few annotated images, requiring the design of label-efficient training solutions [16].
A transfer learning framework, typically fine-tuning ImageNet pre-trained models on medical downstream tasks, has been a common strategy to tackle this challenge [17, 18, 19]. Whether specific in-domain pre-training may benefit the model performance on the downstream task is still debated, and recent works provide controversial results depending on the specific task and domain [20, 9]. Nevertheless, involving in-domain data may be an effective alternative to ImageNet pre-training, especially in a self-supervised fashion to overcome the limited availability of annotations.
Recent advances in self-supervised learning (SSL) methods such as MoCoV3 [21], SimCLRV2 [22], and DINO [23] have shown significant potential. These approaches enable models to learn robust representations from unlabeled data, thereby reducing the dependency on large labeled training datasets. In this context, our main contribution is the proposal of a novel few-shot self-supervised pre-training approach, based on denoising diffusion probabilistic models (DDPMs), specifically focusing on landmark detection for x-ray images, widely used for anatomical assessment and diagnosis, in a realistic and typical scenario where the number of available annotations and images is minimal (up to $50$).
@@
To our knowledge, this work is the first to employ diffusion models for self-supervision in landmark detection. We compare the proposed method with state-of-the-art alternatives including ImageNet supervised pre-training, MoCoV3, SimCLRV2, and DINO self-supervised pre-training, in terms of performance and robustness at the varying of labeled training available samples. Our results show that our method outperforms ImageNet supervised pre-training and the other self-supervised methods for the landmark detection task. Furthermore, we perform a set of experiments to provide insights on whether the usage of a different in-domain pre-training dataset may benefit the downstream task, in a few shot regimes.
@@
The remainder of the paper is organized as follows: Section 2 reviews related studies, particularly focusing on self-supervised pre-training in the medical domain and landmark detection. Section 3 describes our proposed approach. Section 4 gives an overview of the benchmark datasets employed in this research, details the implementation specifics, and explains the evaluation metrics. Lastly, Section 5 showcases our experimental results and insights, drawing some conclusions in Section 6.
@@ 2 Related Works
The problem of few-shot learning is particularly relevant in the medical application domain, as obtaining both images and high-quality, unbiased annotations is costly and time-consuming. Consequently, several works leverage self-supervised pre-training for label efficiency [24, 25, 26, 27].
Zhu et al.[27] propose UOD (Universal One-shot Detection), a domain-adaptive framework for one-shot anatomical landmark detection across multiple medical imaging domains. UOD uses contrastive learning and a domain-adaptive transformer, while reducing annotation burden. Rousseau et al. [24] introduce a pre-training method using diffusion models for dental radiography segmentation. Baranchuk et al. [25] use diffusion models to capture high-level semantic information for semantic segmentation. Brempong et al. [26] propose Decoder Denoising Pretraining (DDeP) to enhance label efficiency in semantic segmentation models by emphasizing the importance of pre-training both encoder and decoder components with limited labeled examples.
In the specific context of landmark detection, the existing works have mainly tackled the data scarcity issue by leveraging transfer learning techniques [28, 8, 9]. Tiulpin et al. [28] use transfer learning from low-budget annotations for knee x-ray landmark localization in osteoarthritis stages. Di Via et al. [9] study whether in-domain data provide benefits over ImageNet pre-training for landmark detection in x-ray images. Zhu et al. [8] propose GU2Net, a universal model for anatomical landmark detection in medical images, addressing limitations in specialized, dataset-dependent methods. Unlike the previously cited works, in this paper, we examine the impact of self-supervised pre-training with diffusion models on landmark detection for x-ray images, focusing on a realistic and typical scenario where the number of available annotations and images is limited. To our knowledge, this is the first application of diffusion models for SSL in this specific task.
@@ 3 Approach
Fig. 1 illustrates our two-step methodology: initially, a DDPM is pre-trained on the training set, in a self-supervised fashion (with no annotations). Subsequently, this model is fine-tuned on the labeled training set for the downstream task, represented by the landmark detection. Further details on these steps are discussed in the following section.
@@
Pre-training using the DDPM: Denoising Diffusion Probabilistic Models [29] (DDPMs) have emerged as a powerful class of generative models, excelling in the synthesis of high-fidelity samples in diverse fields, including medical imaging [30, 31]. Their training methodology involves a forward diffusion process that gradually transforms data into noise and a reverse process that learns to reverse this transformation, thereby generating new data samples from noise.
In this work, we leverage a DDPM Unet architecture as a predictive model for noise. During pre-training, the DDPM Unet receives batches of images alongside their respective diffusion time steps, and it is tasked with predicting the perturbation in the data. Real-world landmark detection datasets are typically small-scale, thus, we expect our DPPM Unet to be pre-trained with small-scale unannotated datasets, while still being able to provide rich and general features for the downstream task. This allows us to keep computational resources and time under a budget, while providing an effective few-shot learning approach for landmark detection.
Fine-tuning on the downstream task: The pre-trained Unet model integrated into the DDPM undergoes fine-tuning for landmark detection in a few-shot regime, thus using a few available annotated data for the downstream task. This step requires only modifying the final classification layer to align with the number of landmarks to be predicted. Although the first stage adopts a self-supervised strategy (with no annotations), the second phase employs a supervised approach by providing the model with ground-truth heat maps for each landmark to be predicted.
@@ 4 Experimental Setup
In this section, we describe the datasets used, the metrics employed, and the details of the implementation of our DDPM architectures and training procedures for the different pre-training methods discussed above. Afterwards, we show the results that have been achieved by evaluating the approach presented.
4.1 Datasets Overview
This section outlines the datasets employed in our study, including Chest x-rays, Cephalometric x-rays, and Hand x-rays. We describe the dataset composition, annotation protocols, and the division into training, validation, and testing sets to ensure clarity and reproducibility in our experiments.
Chest dataset: This dataset comprises $279$ Chest x-ray images sourced from the China set, as documented in [32], and is accessible through the U.S. National Library of Medicine’s public repository. The original image dimensions are approximately $3000$x3000300030003000 pixels. Following the procedures provided in [8, 9], we split the dataset with the first $195$ images for training, the subsequent $34$ for validation and the final $50$ for testing. We evaluate our model’s performance by pixel distance using $6$ manually annotated landmarks per image as all the PNG dataset has no information on physical spacing.
Cephalometric dataset: The Cephalometric dataset [33] is an open-source resource containing $400$ cephalometric x-ray images of individuals aged between $7$ and $76$ years. Each image has a resolution of $0.1$mmx0.10.10.10.1mm and dimensions of $2400$x1935193519351935 pixels. In our research, the first $150$ images are used for training purposes, while the remaining $250$ images are set aside for testing, following the methodology in [8]. To ensure accurate anatomical landmark annotations, two orthodontists identified $19$ landmarks on each image, and the mean coordinates were calculated to address inter-observer variability.
Hand dataset: This open dataset entails $909$ Hand x-ray images, each at an average size of $1563$x2169216921692169 pixels. Consistent with the approaches taken by [8, 9], we allocate the first $550$ images for training, the subsequent $59$ for validation, and the final $300$ images for testing. To enable a direct comparison with other approaches in terms of physical distances we take the estimate made by Payer et al.[34] that the width of the wrist is $50$mm. Then, physical distance is computed as $\frac{50}{|p-q|_{2}}$, where $p$ and $q$ are the two ends of the wrist, determined by first and fifth landmarks in the set of $37$ manually labelled by Payer et al. [34] as well.
(a) Chest landmarks: 6
(b) Cephalometric landmarks: 19
(c) Hand landmarks: 37
Figure 2: Annotated landmarks in x-ray images for Chest, Cephalometric, and Hand datasets.
@@ 4.2 Implementation Details
This section provides the technical specifics of the implementation and experimental setup. The experiments were conducted on a single NVIDIA A30 GPU with 24 GB of RAM, using Python 3.10 and PyTorch 2.1.0.
Data pre-processing: Images are resized to $256$x256256256256 pixels and normalized to a $[0-1]$ range. Post-resizing, random augmentations are applied during training, including rotations (uniformly between $-2$°and $2$°), scaling (from $-0.02$ to $0.02$), and translations (from $-0.02$ to $0.02$).
Pre-training stage: The DDPM is trained with a batch size of $4$, with gradient accumulation every $8$ batches. Training involves $500$ diffusion steps with a linear noise schedule. The training process spans $10k$ iterations, and we save the weights at iterations $4k$, $6k$, $8k$, and $10k$ for comparison during fine-tuning. We utilize the AdamW optimizer with a learning rate of $1e^{-4}$ and apply the exponential moving average to model parameters with a decay factor of $0.995$. We employ mean squared error as the loss function, consistent with the paper [29].
@@
Concerning the implementation of other SSL and ImageNet pre-training techniques, the sources used are [35] and [36].
Fine-tuning stage: The fine-tuning process is carried out using the AdamW optimizer for $200$ epochs, with an early stopping criterion. The learning rate is adjusted using a ReduceLROnPlateau scheduler based on the validation loss, starting at an initial value of $1e^{-5}$. The batch size is set at $2$, with a gradient accumulation of $8$. We use cross-entropy loss for our semantic landmark detection task to predict different classes.
4.3 Evaluation Metrics
In our research, we adopt the same evaluation metrics chosen by [8, 28, 9, 7], namely Mean Radial Error (MRE) and Success Detection Rate (SDR), which are widely used to assess the performance of landmark detection algorithms.
The MRE measures the accuracy of predicted landmarks by calculating the average Euclidean distance to the corresponding ground truth landmarks. For $n$ landmarks, the MRE is defined as:
$MRE=\frac{1}{n}\sum_{i=1}^{n}\sqrt{(x_{i}-\hat{x}_{i})^{2}+(y_{i}-\hat{y}_{i})^{2}}$
where $(x_{i},y_{i})$ represent the ground truth landmark coordinates and $(\hat{x}_{i},\hat{y}_{i})$ represent the predicted coordinates.
SDR, in contrast, evaluates robustness by indicating the percentage of predicted landmarks that fall within a certain threshold distance from the actual landmarks. For $n$ landmarks with a threshold $t$, the SDR is calculated as:
$SDR=\frac{1}{n}\sum_{i=1}^{n}I(d_{i}\leq t)$
where $d_{i}$ is the Euclidean distance between the ground truth and the predicted landmark, and $I$ is an indicator function that returns $1$ if $d_{i}\leq t$ and 00 otherwise.
5 Results and Analysis Discussion
In this section, we present the experiments conducted to evaluate our proposed methods and the corresponding results. We detail the tuning of DDPM pre-training iterations, assess the performance for landmark detection with few labeled data, and analyze the impact of different pre-training datasets on the downstream task.

### v3

@@ Abstract
Deep neural networks have been extensively applied in the medical domain for various tasks, including image classification, segmentation, and landmark detection. However, their application is often hindered by data scarcity, both in terms of available annotations and images. This study introduces a novel application of denoising diffusion probabilistic models (DDPMs) to the landmark detection task, specifically addressing the challenge of limited annotated data in x-ray imaging. Our key innovation lies in leveraging DDPMs for self-supervised pre-training in landmark detection, a previously unexplored approach in this domain. This method enables accurate landmark detection with minimal annotated training data (as few as $50$ images), surpassing both ImageNet supervised pre-training and traditional self-supervised techniques across three popular x-ray benchmark datasets. To our knowledge, this work represents the first application of diffusion models for self-supervised learning in landmark detection, which may offer a valuable pre-training approach in few-shot regimes, for mitigating data scarcity. To bolster further development and reproducibility, we provide open access to our code and pre-trained models for a variety of x-ray related applications: https://github.com/Malga-Vision/DiffusionXray-FewShot-LandmarkDetection
@@ 1 Introduction
Landmark detection, the task of identifying anatomical keypoints in images [14], plays a crucial role in various medical applications, including angle measurements [21], skeletal assessments [19], and surgical planning [8]. Although deep neural networks have become prevalent in this domain [29, 22], and many fully supervised methods have been proposed [43, 44, 33, 15, 36, 9, 39, 34, 16], their effectiveness is often limited by the scarcity of annotated data, particularly in medical imaging where expert annotations are costly and time consuming [44, 43, 33]. Consequently, real-world datasets typically contain very few annotated images, requiring the design of label-efficient training solutions [42].
To address this challenge, researchers have explored transfer learning approaches, typically fine-tuning ImageNet pre-trained models on medical tasks [18, 1, 38]. However, the efficacy of in-domain pre-training remains a subject of debate, with recent studies yielding conflicting results across different tasks and domains [31, 33]. Self-supervised learning (SSL) methods such as MoCoV3 [6], SimCLRV2 [5], and DINO [4] have emerged as promising alternatives, enabling models to learn robust representations from unlabeled data. These approaches have shown potential to reduce the dependency on large labeled datasets, especially for medical imaging applications.
In this context, our primary contribution lies in the novel application of DDPM to the task of landmark detection in x-ray images. While DDPMs have demonstrated remarkable success in image generation tasks [7, 24], their potential for self-supervised pre-training in medical image analysis, particularly landmark detection, remains largely unexplored. We propose a few-shot self-supervised pre-training approach specifically tailored for landmark detection in x-ray images, widely used for anatomical assessment and diagnosis, addressing the realistic and typical scenario where the available annotations and images are minimal (up to $50$).
To our knowledge, this work represents the first application of diffusion models for self-supervision in landmark detection. We conduct a comprehensive comparison of our method against state-of-the-art alternatives, including YOLO [44], ImageNet supervised pre-training and self-supervised approaches like MoCoV3, SimCLRV2, and DINO. Our evaluation focuses on performance and robustness across varying quantities of labeled training samples.
@@
Our results demonstrate that our DDPM-based approach outperforms both ImageNet supervised pre-training and other self-supervised methods in the landmark detection task. Furthermore, we perform a set of experiments to provide insights on whether the usage of a different in-domain pre-training dataset may benefit the downstream task, in a few shot regimes.
To facilitate further research in x-ray imaging, we have made our code and pre-trained models (including DDPM, MoCoV3, SimCLRV2, and DINO) publicly available. These resources, trained on the x-ray datasets used in this study, can be accessed at https://github.com/Malga-Vision/DiffusionXray-FewShot-LandmarkDetection.
This work not only introduces a novel application of DDPMs for the landmark detection task but also contributes to the ongoing discussion on effective pre-training strategies for medical imaging tasks with limited annotations.
@@
The remainder of the paper is organized as follows: Section 2 reviews related studies, particularly focusing on self-supervised pre-training in the medical domain and landmark detection. Section 3 describes our proposed approach. Section 4 gives an overview of the benchmark datasets employed in this research and explains the evaluation metrics. Lastly, Section 5 showcases our experimental results and insights, drawing some conclusions in Section 6.
@@ 2 Related Works
Few-shot learning is particularly relevant in medical applications due to the high cost and time-intensive nature of obtaining high-quality, unbiased annotations [27]. To address this challenge, researchers have explored various approaches, with transfer learning and self-supervised learning emerging as two prominent strategies.
In landmark detection, existing approaches have primarily used transfer learning to address data scarcity [30, 44, 33]. Tiulpin et al. [30] applied transfer learning from low-budget annotations for knee x-ray landmark localization in osteoarthritis stages. Di Via et al. [33] investigated the potential benefits of in-domain data over ImageNet pre-training for x-ray image landmark detection. Zhu et al. [44] proposed GU2Net, a universal model for anatomical landmark detection in medical images, addressing limitations in dataset-dependent methods.
While transfer learning has shown promise, recent advancements in self-supervised learning have opened new avenues for tackling data scarcity in medical imaging tasks [32]. Yao et al. [40] developed a two-stage framework called Cascade Comparing to Detect (CC2D) that uses self-supervised learning and pseudo-label training. Zhu et al. [42] introduced UOD, a domain-adaptive framework using contrastive learning and a domain-adaptive transformer. Miao et al. [23] proposed FM-OSD, using pre-trained visual foundation models for anatomical landmark detection introducing distance-aware similarity learning loss, and bidirectional matching strategy to enhance detection accuracy.
In parallel with these developments, diffusion models have gained traction in various medical imaging tasks [17], although their potential for landmark detection remains largely unexplored. Pinaya et al. [26] developed a latent diffusion model for for 3D brain MRI generation. Rousseau et al. [28] used diffusion models for dental radiography segmentation. Baranchuk et al. [2] employed diffusion models for semantic segmentation. Wyatt et al. [37] proposed AnoDDPM for anomaly detection in medical images.
Building upon these advancements in both self-supervised learning and diffusion models, our work distinguishes itself by examining the impact of self-supervised pre-training with diffusion models on landmark detection for x-ray images in a realistic scenario with limited annotations and images, a common challenge in medical imaging.
To our knowledge, this study represents the first application of diffusion models for self-supervised learning in the specific task of landmark detection in x-ray images. This novel approach bridges the gap between the established potential of diffusion models in other medical imaging tasks and the pressing need for label-efficient methods in landmark detection.
@@ 3 Approach
Fig. 1 illustrates our two-step methodology: initially, a DDPM is pre-trained on the training set, in a self-supervised fashion (with no annotations). Subsequently, this model is fine-tuned on the labeled training set for the downstream task, represented by the landmark detection. Further details on these steps are discussed below.
@@
DDPMs background. Denoising Diffusion Probabilistic Models (DDPMs) [12] have emerged as a powerful class of generative models, excelling in the synthesis of high-fidelity samples in diverse fields, including medical imaging [7, 24]. DDPMs comprise two key processes: forward diffusion and reverse diffusion.
1. Forward Diffusion Process:
The forward diffusion process is described by:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ (1)
Here, $q(x_{t}|x_{t-1})$ represents the probability distribution of $x_{t}$ given $x_{t-1}$. $\mathcal{N}$ denotes a Gaussian distribution with mean $\sqrt{1-\beta_{t}}x_{t-1}$ and variance $\beta_{t}\mathbf{I}$. $\beta_{t}$ is a variance schedule that increases with $t$, controlling the amount of noise added at each step. $\mathbf{I}$ is the identity matrix. The process can be expressed as a single step:
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (2)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. This equation allows direct sampling of $x_{t}$ given $x_{0}$, with $\sqrt{\bar{\alpha}_{t}}x_{0}$ as the mean and $(1-\bar{\alpha}_{t})\mathbf{I}$ as the variance.
2. Reverse Diffusion Process:
The reverse diffusion process is modeled as:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ (3)
Here, $p_{\theta}(x_{t-1}|x_{t})$ is the probability of $x_{t-1}$ given $x_{t}$. $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the mean and covariance predicted by a neural network with parameters $\theta$. The mean of the reverse process is estimated by:
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right)$ (4)
where $\epsilon_{\theta}(x_{t},t)$ is the noise predicted by the neural network. The terms $\frac{1}{\sqrt{\alpha_{t}}}$ and $\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}$ are scaling factors that adjust for the noise level at time t.
The training objective is:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right]$ (5)
This loss function trains the model to predict the noise $\epsilon$ added at each timestep. $\epsilon_{\theta}(\cdot)$ is the noise predicted by the model, and $\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ is the noisy input to the model at time t. The expectation $\mathbb{E}_{t,x_{0},\epsilon}$ is taken over random timesteps, images, and noise samples.
Pre-training using the DDPM. In this work, we leverage a DDPM U-Net architecture as a pre-training strategy for anatomical landmark detection using unannotated medical images. This approach captures multi-scale anatomical representations, with DDPM’s denoising process extracting features at various levels and U-Net’s skip connections preserving both fine details and global information. During pre-training, the model learns to generate new images with the same pixel distribution as the originals, enabling it to learn important anatomical features.
Given the typical scarcity of labeled data in real-world landmark detection, our DDPM U-Net is designed for effective pre-training on small-scale unannotated datasets while maintaining computational efficiency and minimizing training time. This approach yields rich and general features for the downstream task, offering an effective few-shot learning framework for landmark detection. By leveraging unannotated medical data, our model develops a comprehensive understanding of anatomical structures, creating a robust foundation that can adapt to specific landmark detection tasks with minimal additional training.
Ground-truth heatmaps generation. We formulate landmark detection as a classification problem over all image pixels to generate output heatmaps. This approach enables simultaneous prediction of multiple landmarks, with one heatmap per landmark. Specifically, our model outputs $N$ heatmaps, where $N$ is the number of landmarks to predict. For heatmap generation, we adopt a Gaussian-based approach inspired by [44]. This strategy involves applying a Gaussian filter ($\sigma=5$) centered at the ground-truth landmark location, followed by thresholding to obtain binary images. Formally:
$H_{g}(i,j)=\begin{cases}1,&\text{if }\exp\left(-\frac{(i-x)^{2}+(j-y)^{2}}{2\sigma^{2}}\right)>\frac{1}{2}\max(G)\\ 0,&\text{otherwise}\end{cases}$
Where $H_{g}(i,j)$ is the Gaussian heatmap pixel value at $(i,j)$, $(x,y)$ are the ground truth landmark coordinates, and $G$ is the Gaussian-filtered image before thresholding.
Training on these heatmaps enables our model to learn landmark spatial relationships, predicting both general areas of interest and precise coordinates, and potentially enhance downstream landmark detection.
Fine-tuning on the downstream task. The pre-trained DDPM UNet model is adapted for anatomical landmark detection, bridging the gap between generative modeling and precise anatomical localization. This adaptation, illustrated in Figure 1, involves a subtle yet powerful modification of the UNet architecture. While preserving the core structure that proved so effective in the DDPM, we alter the final convolutional layer to output multiple channels (each corresponding to a predicted heatmap for a specific anatomical landmark) and the timestep embedding is set to be null (as the temporal aspect is irrelevant for static landmark detection). This architectural shift transforms the model from a general image generator to a specialized landmark detector, capitalizing on learned hierarchical representations.
Our fine-tuning strategy transitions from DDPM’s self-supervised paradigm to a supervised approach tailored for landmark detection. Training with ground-truth heatmaps allows the model to repurpose its learned features for precise spatial localization. Landmark coordinates are derived by calculating the centroid of each heatmap generated.
This supervised fine-tuning, combined with pre-trained weight transfer (indicated by the blue dashed arrow), enables quick adaptation to the new task with minimal annotated data. The approach’s efficacy in few-shot learning contexts highlights its potential for medical imaging applications, where large annotated datasets are often scarce. By harnessing the power of diffusion models, traditionally used for generation tasks, we show their unexplored potential in discriminative tasks like landmark detection.
@@ 4 Experimental Setup
4.1 Datasets and Evaluation Metrics
We evaluate our method on three public x-ray datasets standard in landmark detection research, following the dataset splits and evaluation protocols of [44, 33]. The Chest x-ray dataset [3] comprises $279$ images (approximately $3000$×$3000$ pixels, resolution unknown) with $6$ landmarks, split into $195$ for training, $34$ for validation, and $50$ for testing. The Cephalometric x-ray dataset [35] consists of $400$ lateral cephalograms ($2400$×$1935$ pixels, $0.1$ mm resolution) with $19$ target landmarks, split into $130$ for training, $20$ for validation, and $250$ for testing. The Hand x-ray dataset [10] contains $909$ radiographs (average size $1563$×$2169$ pixels) with $37$ labeled landmarks, assuming a $50$ mm length between wrist endpoints [25], split into $550$ for training, $59$ for validation, and $300$ for testing.
Our model is assessed using two standard metrics [27, 41]: the mean radial error (MRE) and the successful detection rate (SDR). MRE measures the Euclidean distance between predicted and ground truth landmarks, while SDR calculates the percentage of predictions within dataset-specific thresholds. For the Chest dataset, we report these metrics in pixels due to unavailable physical spacing information, whereas for Cephalometric and Hand datasets, we use millimeters, adhering to field standards.
@@ 4.2 Implementation Details
The experiments are conducted on a single NVIDIA A30 GPU with 24 GB RAM, using Python 3.10 and PyTorch 2.1.0. For data pre-processing, images are resized to $256$x256256256256 pixels, normalized to $[0-1]$ range, and augmented during training with random rotations ($-2$°and $2$°), scaling ($-0.02$ to $0.02$), and translations ($-0.02$ to $0.02$).
Pre-training stage. The DDPM is trained with a batch size of $4$, with gradient accumulation every $8$ batches. Training involves $500$ diffusion steps with a linear noise schedule. The training process spans $10k$ iterations, and we save the weights at iterations $4k$, $6k$, $8k$, and $10k$ for comparison during fine-tuning. We utilize the AdamW optimizer with a learning rate of $1e^{-4}$ and apply the exponential moving average to model parameters with a decay factor of $0.995$. We employ mean squared error as the loss function, consistent with the paper [12].
@@
Concerning the implementation of other SSL and ImageNet pre-training techniques, the sources are [11] and [13].
Fine-tuning stage. The fine-tuning process is carried out using the AdamW optimizer for $200$ epochs, with an early stopping criterion. The learning rate is adjusted using a ReduceLROnPlateau scheduler based on the validation loss, starting at an initial value of $1e^{-5}$. The batch size is set at $2$, with a gradient accumulation of $8$. We use cross-entropy loss for gaussian heatmaps, and a negative log-likelihood loss for the contour-hugging ones [20].
5 Experimental Results